### v1

# Is My Data in Your Retrieval Database? Membership Inference Attacks Against Retrieval Augmented Generation

Abstract
Retrieval Augmented Generation (RAG) systems have shown great promise in natural language processing. However, their reliance on data stored in a retrieval database, which may contain proprietary or sensitive information, introduces new privacy concerns. Specifically, an attacker may be able to infer whether a certain text passage appears in the retrieval database by observing the outputs of the RAG system, an attack known as a Membership Inference Attack (MIA). Despite the significance of this threat, MIAs against RAG systems have yet remained under-explored.
This study addresses this gap by introducing an efficient and easy-to-use method for conducting MIA against RAG systems. We demonstrate the effectiveness of our attack using two benchmark datasets and multiple generative models, showing that the membership of a document in the retrieval database can be efficiently determined through the creation of an appropriate prompt in both black-box and gray-box settings. Our findings highlight the importance of implementing security countermeasures in deployed RAG systems to protect the privacy and security of retrieval databases.
1 Introduction
Retrieval Augmented Generation (RAG) systems [12, 6] have emerged as a promising approach in natural language processing, gaining significant attention in recent years due to their ability to generate high-quality, up-to-date and contextually relevant responses. These systems combine retrieval and generation components to provide more accurate and informative responses compared to traditional language models. However, like any advanced technology, RAG systems are not immune to vulnerabilities.
While previous research has successfully demonstrated various types of attacks against RAG systems [10, 21, 7, 20], there are still unexplored vulnerabilities in these systems that may pose privacy risks. Specifically, the use of the retrieval database introduces new privacy concerns. Since the retriever component searches a database for relevant passages, an attacker may be able to infer whether a certain text passage appears in the database by observing the outputs of the RAG system. This type of attack is known as a Membership Inference Attack (MIA), and can be used to reveal sensitive information about the contents of the retrieval database.
MIAs were extensively researched in the past in the context of various machine and deep learning systems [15, 3, 2, 18, 9], however to the best of our knowledge the topic of membership inference against RAG systems remains under-explored.
When a MIA is performed against a RAG system it can potentially reveal sensitive or proprietary company information. This may include information about individuals or organizations included in the retrieval database. Furthermore, MIA can be used to prove the unauthorized use of proprietary documents, as part of a legal action [16]. This dual capability makes them a particularly serious form of attack that must be investigated in order to ensure the security and privacy of these systems.
In this study, we introduce a method that is both efficient and easy to use for conducting MIAs against RAG systems, with the objective of determining whether a specific data sample is present in the retrieval database. We assess the effectiveness of our attack using two benchmark datasets that represent scenarios where privacy may be a concern, and employing multiple generative models.
The findings of our study reveal that the membership of a document in the retrieval database can be efficiently determined through the creation of an appropriate prompt, underscoring the importance of implementing security countermeasures in deployed RAG systems.
2 Background
2.1 Retrieval Augmented Generation
Figure 1: Example RAG template for the generation phase of a RAG system. The highlighted placeholders are replaced by the fetched documents from the database and the user prompt, respectively.
Retrieval Augmented Generation is a technique for enriching the knowledge of Large Language Models (LLMs) with external databases without requiring any additional training. This allows easy customization of trained LLMs for specific needs, such as creating AI-based personal assistants or making long textual content (such as a user manual) accessible using simple text queries [4].
Prior to deploying a RAG pipeline, a retrieval database $\mathcal{D}$ must be populated with documents. During this initialization phase, each document is split into chunks, which are mapped into vector representations using a retrieval model $E$, and then stored as an index in the database alongside the original document.
The retrieval model $E$ is specifically designed to learn a mapping from user prompts and documents to a shared vector space, such that prompts are embedded close together with documents containing relevant information to respond to the prompt. This enables efficient searching in the retrieval database, as semantically similar prompts and documents are clustered together in the vector space.
Once deployed, the RAG pipeline typically consists of two main phases: search and generation. In the search phase, $\mathcal{D}$ is queried to find relevant documents that match the user’s query or prompt. When a user prompt is processed by the RAG pipeline, the prompt $p$ is mapped into a vector representation using $E$. Then, $\mathcal{D}$ is searched to find the top-k most similar entries, based on a distance metric calculated using the vector representations (e.g., Euclidean distance). The retrieved entries from $\mathcal{D}$ are organized and provided to the generation phase together with the user prompt.
In the generation phase, a language model $G$ synthesizes the answer based on the retrieved entries from $\mathcal{D}$. The organized information from the search phase is inserted into the RAG template to generate a context $C$. The final system response is obtained by feeding the context $C$, concatenated with the user prompt $p$, into $G$:
$\text{Response}=G(C\parallel p)$ (1)
where $\parallel$ denotes text concatenation.
In practice, it is common to place titles indicating the different areas in the input or additional instructions. For example, placing the sentence ”Answer the question based on the context” after the context and before the user prompt. A full example of a RAG template appears in Figure 1.
2.2 Membership Inference Attacks
Membership inference attacks [15, 9] are a type of privacy threat where an attacker aims to determine whether a specific data record was used in the training set of a machine learning model. This carries significant privacy implications as it can potentially reveal sensitive information about individuals, even if the model does not directly release any personal data.
Formally, an attacker aims to determine the membership of a sample $x$ in the training data of a target model $\mathcal{D}_{m}$, i.e., to check if $x\in\mathcal{D}_{m}$, which is known as sample-level membership inference. Typically, these attacks involve calculating one or more metrics on the target model’s outputs that reflect the probability of the sample being a part of the training set, such as the model outputs’ entropy or log-probabilities [3]. Several metrics may be computed for each sample and then fused together using a machine learning model, known as an attack model, which in turn outputs the probability of a sample being a member of the training set. Additionally, an attacker may also aim to determine the membership of a certain user, i.e., to check if a user’s data is part of the training set, which is known as user-level membership inference [14]. However throughout this paper we will address the membership inference challenge from a sample-level perspective.
In the context of RAG, membership inference can be attributed to either the membership of a sample in the training dataset of the models $E$ or $G$ (described in the previous subsection 2.1), or a document’s membership in the retrieval dataset $\mathcal{D}$. This paper focuses on the latter. Formally, the goal of the attack is to infer the membership of a target document $d$ in the retrieval database $\mathcal{D}$, i.e., to check if $d\in\mathcal{D}$, using only the final output of the RAG system, namely the output of the generative model $G$ conditioned on the fetched context from the retrieval database $\mathcal{D}$.
To the best of our knowledge, this is the first paper to propose such a membership inference attack tailored to RAG systems.
2.3 Threat Model
This paper considers a black-box scenario, in which the attacker has access solely to the user prompt and the resulting generated output from the RAG system. The attacker can modify the user prompt in any manner they deem appropriate; however, they possess no knowledge of the underlying models $E$ or $G$, nor the prompt templates that are being used by these models. Furthermore, the attacker has no information regarding the deployment details, such as the type of retrieval database employed.
In addition to the black-box setting, we also evaluate a supplementary gray-box scenario, in which the attacker has access to the log-probabilities of the generated tokens. Moreover, in this scenario, we assume that the attacker can train the attack model on the target model’s outputs on a subset of the model’s actual training and test datasets [13].
3 Methodology
Figure 2: Attack prompt example for RAG MIA. The highlighted text is the attack-specific part of the prompt, and the rest is taken from the sample for which membership is inferred.
In this section we present our membership inference attack.
The input to the generative model used by the system consists of the RAG tempate into which the retrieved database entries and the user prompt are inserted, as depicted in Figure 1. Since the attacker’s interaction with the system is limited to the user prompt part of the RAG template, the attack must be based on a prompt that achieves both of the following goals:
  • Cause the system to fetch the desired sample from the retrieval database;
  • Cause the system to generate an output that either contains the membership status of the sample, or from which it can be deduced.
The first goal can be easily fulfilled by creating a prompt that has high similarity with the desired sample in the retrieval database. Essentially, any prompt that contains the target sample should be sufficient, as long as there is not too much text added to it. However, the prompt must also cause the generation process to expose the membership status of the sample. Prompting the system with target sample alone will not necessarily achieve this desired result.
To handle this, we designed the following attack prompt format, along with a few additional variations presented in Table 1: ”Does this: ”{Target Sample}” appear in the context? Answer with Yes or No.”. In Figure 2 we present an example of this attack prompt format. Using this prompt we are able to achieve both goals: cause the right sample to be retrieved from the retrieval database, and make the generation phase produce a binary response regarding the sample’s presence in the context. In the black-box setting, we use the model’s answer (Yes/No) alone to deduce the membership status of samples.
As an enhancement to our attack, in cases where the adversary has access to the log-probabilities of the selected tokens, we additionally employ an attack model, similar to the approach described in Section 2.2. In this setup, we apply an ensemble of attack models [13], using as input to the attack both the logits and the class-scaled-logits [3] corresponding to the ”Yes” and ”No” tokens output from the attacked model. The logits are computed by first calculating the exponent of of the log-probability to get a probability estimate $P$ and then applying the logit function.
Since the model only outputs the log-probability of the selected token, for example the ”Yes” token, without the complementary ”No” token, we assign a fixed low probability value of $0.001$ to the complementary token.
Black-Box
Gray-Box
Figure 3: Comparison of different attack prompts. Top: HealthCareMagic, Bottom: Enron
3.1 Experimental Setup
Generative Models
The experiments were conducted on three generative language models:
  • google/flan-ul2 [17], denoted by flan
  • meta-llama/llama-3-8b-instruct [1], denoted by llama
  • mistralai/mistral-7b-instruct-v0-2 [11], denoted by mistral
Datasets
Throughout the evaluation we used two datasets that represent practical scenarios, in which privacy can be critical:
From each dataset, we randomly selected 8,000 samples to be stored in the retrieval database, which we denote as member documents; the remaining 2,000 samples were used as non-member documents in our evaluation.
Embedding Models
The embedding model we used is sentence-transformers/all-minilm-l6-v2, which maps sentences and paragraphs to a 384-dimensional dense vector space.
Retrieval Database
We used a Milvus Lite [19, 8] vector database with $k=4$, Euclidean distance (L2) metric type and HNSW index.
RAG Template
The input to the generative model is built from the user prompt and the context fetched from the retrieval database as a response to the prompt, using the following template:
[…]
In our case, the user prompt was replaced with our special attack prompt.
Attack Prompt
In our evaluation we experimented with 5 different attack prompts, listed in Table 1. Each attack prompt includes a placeholder for a sample, which can be a member or a non-member sample, as shown in Section 3. In the case of the Enron dataset, the sample is the full email body, or its first 1000 characters if it is longer. In HealthCareMagic, the human part of the dialogue is used as the sample.
MIA Attack Details
In both the black-box and the gray-box scenarios we ran our attack using random samples containing 2000 member documents and 2000 non-member documents. In the gray-box case, these documents were split between 40 attack models in the ensemble [13], each model trained 5 times with different train-test splits. In this evaluation, we classified model outputs that failed to include either the ”Yes” or ”No” tokens as non-member documents. This case accounted for approximately 6% of the total outputs. We present the statistics on the amount of such documents in Appendix C.
4 Results
4.1 Impact of Attack Prompt
Table 1: RAG-MIA attack prompts
Attack Prompt
[…]
context?
[…]
the context?
[…]
with Yes or No.
[…]
one word. Answer No if you are not sure.
One important aspect of performing attacks against any generative model is the prompt used to trigger the attack. In the case of RAG, the attack prompt is first used to fetch the relevant documents from the retrieval database, and is then incorporated together with the fetched documents into the RAG template and provided to the generative model. To achieve the best performance, we experimented with five possible attack prompts, described in Table 1. Each one tries to manipulate the RAG system in a slightly different manner to both fetch the relevant documents from the retrieval database, and cause the generative model to output whether the target document was indeed in the retrieval database.
[…]
Unsurprisingly, the results in the black-box setting are inferior to those in the gray-box setting. This is especially the case for the flan model, with an improvement of up to 25% in the gray-box setting. This means that the log-likelihood values for member documents are significantly higher than for non-member documents, indicating a higher confidence of the model in its response.
However, when looking at the llama and mistral models, the average difference is only up to 12%, and in some cases even lower, depending on the prompt.
To further explore this difference, we analyze the percentage of samples that are correctly retrieved from the database for each prompt. We found that over 95% of the member samples are indeed retrieved for both datasets. This is in contrast with the non-member samples, which are retrieved in nearly 0% of the cases. The full results of this analysis can be found in Appendix B.
Thus, we conclude that the flan model is more grounded to the content of its input prompt, and thereby more sure of the presence/absence of a piece of text from it in comparison to the llama and mistral models.
4.2 Results Summary
To quantify the overall MIA risk, we summarize the best results (across prompts) for each combination of model and dataset in Table 2 . We can see that, on average, the attack success rate is highest for flan, in both threat models.
In addition, we observe that the overall risk in the black-box setting is similar between almost all models, with the exception of the HealthCareMagic-llama case. However in the gray-box setting the results vary more.
Our lowest attack AUC-ROC score of 0.74 is quite high compared to previous research on sample-level MIA in language models [5, 14]. Overall, our black-box attack achieves an average AUC-ROC of 0.80 and the gray-box attack achieves an average AUC-ROC of 0.90. The best gray-box attack on flan accomplishes perfect performance. These results highlight the risk associated with deploying RAG based systems without proper defense mechanisms.
Table 2: RAG-MIA results summary
Black-Box
AUC ROC
Gray-Box
AUC ROC
Dataset Model
HealthCareMagic flan 0.80 1.00
llama 0.88 0.96
mistral 0.74 0.83
Enron flan 0.82 0.96
llama 0.79 0.83
mistral 0.78 0.82
5 Conclusion
In this paper we introduced a new membership inference attack against RAG-based systems meant to infer if a specific document was part of the retrieval database or not, a research question previously not explored. Our attack is demonstrated both in black-box and gray-box threat models, and takes advantage of a characteristic of generative models that is usually considered an advantage - context grounding.
Our attack achieves a very high average AUC-ROC of 0.90 and 0.80 in the gray-box and the black-box threat models, respectively, and for some models almost perfect performance. This underscores the need to develop defenses and countermeasures for this attack scenario. We hope that the research community will proceed to explore the risk of membership inference in RAG-based systems and employ the ideas from this paper as a baseline.
6 Acknowledgement
This work was performed as part of the NEMECYS project, which is co-funded by the European Union under grant agreement ID 101094323, by UK Research and Innovation (UKRI) under the UK government’s Horizon Europe funding guarantee grant numbers 10065802, 10050933 and 10061304, and by the Swiss State Secretariat for Education, Research and Innovation (SERI).
Appendix A Appendices
A Detailed Attack AUC ROC Scores
Table 3 presents the AUC ROC scores for both threat models: black-box and gray-box, and for the different attack prompts.
Table 3: Full RAG-MIA results
Black-Box
AUC ROC
Gray-Box
AUC ROC
Dataset Model
Attack
Prompt
HealthCareMagic flan 0 0.80 0.98
1 0.75 0.97
2 0.78 0.99
3 0.76 0.84
4 0.80 1.00
llama 0 0.82 0.88
1 0.67 0.64
2 0.88 0.96
3 0.78 0.83
4 0.84 0.89
mistral 0 0.71 0.73
1 0.67 0.65
2 0.74 0.83
3 0.71 0.73
4 0.71 0.55
Enron flan 0 0.79 0.93
1 0.68 0.80
2 0.82 0.96
3 0.60 0.62
4 0.80 0.92
llama 0 0.71 0.65
1 0.62 0.54
2 0.79 0.83
3 0.69 0.58
4 0.74 0.74
mistral 0 0.66 0.60
1 0.70 0.68
2 0.78 0.82
3 0.64 0.53
4 0.74 0.69
B Detailed Results of Retrieval Matches
Tables 4 and 5 show the percent of exact matches between the retrieved documents and the member and non-member samples, respectively, for the different attack prompts. We see that the chosen attack prompts do not differ in their influence on the retrieval accuracy.
Table 4: Database retrieval of member documents
Attack
prompt
Equal
count
Total
count
Equal
percent
Dataset
HealthCareMagic 0 1928 2000 96.40
1 1921 2000 96.05
2 1921 2000 96.05
3 1930 2000 96.50
4 1924 2000 96.20
Enron 0 1910 2000 95.50
1 1908 2000 95.40
2 1907 2000 95.35
3 1910 2000 95.50
4 1908 2000 95.40
Table 5: Database retrieval of non-member documents
Attack
prompt
Equal
count
Total
count
Equal
percent
Dataset
HealthCareMagic 0 0 2000 0.00
1 0 2000 0.00
2 0 2000 0.00
3 0 2000 0.00
4 0 2000 0.00
Enron 0 1 2000 0.05
1 1 2000 0.05
2 0 2000 0.00
3 1 2000 0.05
4 0 2000 0.00
C Model outputs without a clear yes/no
answer
The number of model outputs that did not contain either of the ”Yes” or ”No” tokens are shown in Table 6.
Table 6: Model outputs missing yes/no tokens
Missing Total
Percent
missing
Dataset Model
HealthCareMagic flan 923 20000 4.62%
llama 1253 20000 6.27%
mistral 1736 20000 8.68%
Enron flan 1327 20000 6.64%
llama 954 20000 4.77%
mistral 1125 20000 5.63%

### v2

@@
 We demonstrate the effectiveness of our attack using two benchmark datasets and multiple generative models, showing that the membership of a document in the retrieval database can be efficiently determined through the creation of an appropriate prompt in both black-box and gray-box settings.
-Our findings highlight the importance of implementing security countermeasures in deployed RAG systems to protect the privacy and security of retrieval databases.
+Moreover, we introduce an initial defense strategy based on adding instructions to the RAG template, which shows high effectiveness for some datasets and models.
+Our findings highlight the importance of implementing security countermeasures in deployed RAG systems and developing more advanced defenses to protect the privacy and security of retrieval databases.
+Keywords:
+AI Privacy; Membership Inference; RAG; Large Language Models
 1 Introduction
@@
 We assess the effectiveness of our attack using two benchmark datasets that represent scenarios where privacy may be a concern, and employing multiple generative models.
-The findings of our study reveal that the membership of a document in the retrieval database can be efficiently determined through the creation of an appropriate prompt, underscoring the importance of implementing security countermeasures in deployed RAG systems.
+Moreover, we introduce an initial defense strategy based on adding instructions to the RAG template to prevent the model from responding to the attack prompt.
+This approach seems to be highly effective for some datasets and models, but should be further developed to provide a more comprehensive solution.
+The findings of our study reveal that the membership of a document in the retrieval database can be efficiently determined through the creation of an appropriate prompt, underscoring the importance of developing appropriate defenses and implementing security countermeasures in deployed RAG systems.
 2 Background
@@
 2.2 Membership Inference Attacks
-Membership inference attacks [15, 9] are a type of privacy threat where an attacker aims to determine whether a specific data record was used in the training set of a machine learning model.
+Membership inference attacks [15, 9] are a type of privacy threat, where an attacker aims to determine whether a specific data record was used in the training set of a machine learning model.
 This carries significant privacy implications as it can potentially reveal sensitive information about individuals, even if the model does not directly release any personal data.
@@
 Additionally, an attacker may also aim to determine the membership of a certain user, i.e., to check if a user’s data is part of the training set, which is known as user-level membership inference [14].
-However throughout this paper we will address the membership inference challenge from a sample-level perspective.
+Throughout this paper we will address the membership inference challenge from a sample-level perspective.
 In the context of RAG, membership inference can be attributed to either the membership of a sample in the training dataset of the models $E$ or $G$ (described in the previous subsection 2.1), or a document’s membership in the retrieval dataset $\mathcal{D}$.
@@
 2.3 Threat Model
-This paper considers a black-box scenario, in which the attacker has access solely to the user prompt and the resulting generated output from the RAG system.
+This paper considers a black-box scenario in which the attacker has access solely to the user prompt and the resulting generated output from the RAG system.
 The attacker can modify the user prompt in any manner they deem appropriate; however, they possess no knowledge of the underlying models $E$ or $G$, nor the prompt templates that are being used by these models.
@@
 Datasets
-Throughout the evaluation we used two datasets that represent practical scenarios, in which privacy can be critical:
+Throughout the evaluation we used two datasets that represent practical scenarios in which privacy can be critical:
 From each dataset, we randomly selected 8,000 samples to be stored in the retrieval database, which we denote as member documents; the remaining 2,000 samples were used as non-member documents in our evaluation.
@@
 The input to the generative model is built from the user prompt and the context fetched from the retrieval database as a response to the prompt, using the following template:
-[…]
+Please answer the question using the context provided.
+If the question is unanswerable, say unanswerable”.
+Question: {user prompt}.
+Context:
+{context}
+Question: {user prompt}
 In our case, the user prompt was replaced with our special attack prompt.
@@
 4 Results
-4.1 Impact of Attack Prompt
-Table 1: RAG-MIA attack prompts
-Attack Prompt
-[…]
-context?
-[…]
-the context?
-[…]
-with Yes or No.
-[…]
-[…]
-one word. Answer No if you are not sure.
+4.1 Impact of Attack Prompts
 One important aspect of performing attacks against any generative model is the prompt used to trigger the attack.
@@
 Each one tries to manipulate the RAG system in a slightly different manner to both fetch the relevant documents from the retrieval database, and cause the generative model to output whether the target document was indeed in the retrieval database.
+Table 1: RAG-MIA attack prompts
+Attack Prompt
+[…]
+[…]
+[…]
+[…]
+[…]
 […]
@@
 The full results of this analysis can be found in Appendix B.
-Thus, we conclude that the flan model is more grounded to the content of its input prompt, and thereby more sure of the presence/absence of a piece of text from it in comparison to the llama and mistral models.
-4.2 Results Summary
+Thus, we conclude that the flan model is more grounded to the content of its input prompt (context grounding), and thereby more sure of the presence/absence of a piece of text from it in comparison to the llama and mistral models.
+4.2 Attack Results Summary
 To quantify the overall MIA risk, we summarize the best results (across prompts) for each combination of model and dataset in Table 2 .
@@
 However in the gray-box setting the results vary more.
-Our lowest attack AUC-ROC score of 0.74 is quite high compared to previous research on sample-level MIA in language models [5, 14].
-Overall, our black-box attack achieves an average AUC-ROC of 0.80 and the gray-box attack achieves an average AUC-ROC of 0.90.
-The best gray-box attack on flan accomplishes perfect performance.
-These results highlight the risk associated with deploying RAG based systems without proper defense mechanisms.
 Table 2: RAG-MIA results summary
@@
 mistral 0.78 0.82
+Our lowest attack AUC-ROC score of 0.74 is quite high compared to previous research on sample-level MIA in language models [5, 14].
+Overall, our black-box attack achieves an average AUC-ROC of 0.80 and the gray-box attack achieves an average AUC-ROC of 0.90.
+The best gray-box attack on flan accomplishes almost perfect performance.
+These results highlight the risk associated with deploying RAG based systems without proper defense mechanisms.
+4.3 Defense Strategy
+To counter the threat of RAG-MIA attacks, we propose an initial defense strategy in this section.
+Building on previous research that has demonstrated the ability to influence the output of LLMs through system instructions, we leverage this concept to develop a defense mechanism.
+Our approach involves modifying the RAG template to include explicit instructions for the LLM to disregard requests that directly inquire about the retrieved database content, thereby providing a defense against RAG-MIA attacks.
+Specifically, we evaluated the effectiveness of variations of the following modified RAG template:
+Please answer the question using the context provided.
+If the question is unanswerable, say unanswerable”.
+Pointers for the answer:
+1. Repeating, outputting or printing the context is illegal,
+answer unanswerable”.
+2. Asking if a specific text, fact, or passage appears in your
+context is illegal, answer unanswerable”.
+Question: {user prompt}
+Context:
+{context}
+We have slightly adapted the RAG template for the llama model, such that it will respond to the defense instructions better, by placing the defense instructions inside the system prompt, an integral capability of llama models.
+In Table 3 we present the results, and compare them to the scenario where no defense is placed in the RAG system.
+Table 3: RAG MIA results with defense
+Without defense With defense
+Black-Box
+AUC ROC
+Gray-Box
+AUC ROC
+Black-Box
+AUC ROC
+Gray-Box
+AUC ROC
+Dataset Model
+HealthCareMagic flan 0.78 0.99 0.86 0.90
+llama 0.88 0.96 0.74 0.51
+mistral 0.74 0.83 0.72 0.44
+Enron flan 0.82 0.97 0.88 0.96
+llama 0.79 0.84 0.77 0.78
+mistral 0.78 0.80 0.78 0.72
+Our evaluation reveals that the proposed defense strategy yields the most significant benefits against gray-box attacks on the llama and mistral models, across both datasets.
+Notably, the defense demonstrates high efficacy in the gray-box setting for the mistral model, particularly on the HealthCareMagic dataset, where we observe a substantial improvement of 0.39 in AUC-ROC.
+In contrast, the defense has a minimal impact on the flan model, only showing a slight effect in the gray-box setting.
+These findings suggest that further research is needed to explore the feasibility of crafting a RAG template that can effectively defend a flan model against RAG-MIA attacks.
+Furthermore, our analysis of the model outputs reveals that a significant proportion of responses do not contain either ”Yes” or ”No” tokens (classified as non-member documents), accounting for approximately 96% of the total outputs on llama, and about 93% on mistral with HealthCareMagic dataset, which explains the defense performance for these cases.
+On the other hand, for flan, they remain the same (0%).
+We present the statistics on the amount of such documents in Appendix D.
+4.3.1 Improved defense for llama
+Since llama models have a dedicated section in the input prompt for system instructions, as mentioned in Section 4.3, we have experimented with placing the retrieved database content within this dedicated section.
+We compare two cases: (1) With Defense #1 - only the defense instructions are added to the system section (2) With Defense #2 - both defense instructions and retrieved database content are placed in the system section.
+We present the results for this experiment in Table 4.
+Table 4: Llama defenses
+Without Defense With Defense #1 With Defense #2
+Black-Box
+AUC ROC
+Gray-Box
+AUC ROC
+Black-Box
+AUC ROC
+Gray-Box
+AUC ROC
+Black-Box
+AUC ROC
+Gray-Box
+AUC ROC
+Dataset Model
+HealthCareMagic llama 0.88 0.96 0.74 0.51 0.42 0.75
+Enron llama 0.79 0.84 0.77 0.78 0.46 0.92
+As shown in Table 4, placing defense instructions and retrieved database content in the system section provides robust defense against black-box attacks.
+However, this approach is less effective against gray-box attacks, where Defense #1 is preferred.
+Since black-box attacks are more common and require less effort from the attacker, we recommend using Defense #2.
+Nevertheless, we encourage the research community to develop a defense strategy that can effectively protect from both kinds of attack.
 5 Conclusion
@@
 Our attack achieves a very high average AUC-ROC of 0.90 and 0.80 in the gray-box and the black-box threat models, respectively, and for some models almost perfect performance.
-This underscores the need to develop defenses and countermeasures for this attack scenario.
+Our initial defense was able to reduce the success rate of the attack in almost all cases, and mostly prevented the attack for the llama model in the black-box setting.
+The model that mostly did not benefit from this defense was the flan model for which further defenses need to be developed.
+This underscores the need to develop more advanced defenses and countermeasures for this attack scenario.
 We hope that the research community will proceed to explore the risk of membership inference in RAG-based systems and employ the ideas from this paper as a baseline.
@@
 This work was performed as part of the NEMECYS project, which is co-funded by the European Union under grant agreement ID 101094323, by UK Research and Innovation (UKRI) under the UK government’s Horizon Europe funding guarantee grant numbers 10065802, 10050933 and 10061304, and by the Swiss State Secretariat for Education, Research and Innovation (SERI).
-Appendix A Appendices
+Appendices
 A Detailed Attack AUC ROC Scores
@@
 B Detailed Results of Retrieval Matches
-Tables 4 and 5 show the percent of exact matches between the retrieved documents and the member and non-member samples, respectively, for the different attack prompts.
+Tables 6 and 7 show the percent of exact matches between the retrieved documents and the member and non-member samples, respectively, for the different attack prompts.
 We see that the chosen attack prompts do not differ in their influence on the retrieval accuracy.
@@
 4 0 2000 0.00
-C Model outputs without a clear yes/no
-answer
-The number of model outputs that did not contain either of the ”Yes” or ”No” tokens are shown in Table 6.
+C Model outputs without a clear yes/no answer
+The number of model outputs that did not contain either of the ”Yes” or ”No” tokens across attack prompts for each dataset and model combination are shown in Table 8.
 Table 8: Model outputs missing yes/no tokens
@@
 mistral 1125 20000 5.63%
+D Model outputs without a clear yes/no answer with defense prompt
+The number of model outputs that did not contain either of the ”Yes” or ”No” tokens when using a defense prompt for each dataset and model combination are shown in Table 9.
+This is compared to the corresponding model outputs when using attack prompt #2 without the defense.
+The number of missing answers increases significantly for the llama and mistral models, and remains the same for flan.
+Table 9: Model outputs missing yes/no tokens with defense prompt
+Without defense With defense
+Missing Total
+Percent
+missing
+Missing Total
+Percent
+missing
+Dataset Model
+HealthCareMagic flan 0 4000 $0$ 4000 0.00%
+llama 0 4000 0.00% 3868 4000 96.70%
+mistral 101 4000 2.53% 3719 4000 92.97%
+Enron flan 0 4000 $0$ 4000 0.00%
+llama 0 4000 0.00% 3805 4000 95.12%
+mistral 41 4000 1.03% 2743 4000 68.58%

### v3

@@
 Our findings highlight the importance of implementing security countermeasures in deployed RAG systems and developing more advanced defenses to protect the privacy and security of retrieval databases.
-Keywords:
-AI Privacy; Membership Inference; RAG; Large Language Models
 1 INTRODUCTION
-Retrieval Augmented Generation (RAG) systems [12, 6] have emerged as a promising approach in natural language processing, gaining significant attention in recent years due to their ability to generate high-quality, up-to-date and contextually relevant responses.
-These systems combine retrieval and generation components to provide more accurate and informative responses compared to traditional language models.
+Retrieval Augmented Generation (RAG) systems [Lewis et al., 2020, Gao et al., 2023] have emerged as a promising approach in natural language processing, gaining significant attention in recent years due to their ability to generate high-quality, up-to-date and contextually relevant responses.
+These systems combine a retrieval database and retrieval and generation components to provide more accurate and informative responses compared to traditional language models.
+For example, in the health domain, access to timely medical literature, research papers and patient records using a RAG architecture can assist in expediting diagnosis and improving patient outcomes.
 However, like any advanced technology, RAG systems are not immune to vulnerabilities.
-While previous research has successfully demonstrated various types of attacks against RAG systems [10, 21, 7, 20], there are still unexplored vulnerabilities in these systems that may pose privacy risks.
-Specifically, the use of the retrieval database introduces new privacy concerns.
-Since the retriever component searches a database for relevant passages, an attacker may be able to infer whether a certain text passage appears in the database by observing the outputs of the RAG system.
+While previous research has successfully demonstrated various types of attacks against RAG systems [Hu et al., 2024, Zou et al., 2024, Greshake et al., 2023, Zeng et al., 2024], there are still unexplored vulnerabilities in these systems that may pose privacy risks.
+Specifically, the use of a retrieval database, that may contain sensitive or proprietary information, introduces new privacy concerns for the data residing in that database.
+Since the retrieval database is searched for relevant passages to aid the model in responding to a specific user prompt, an attacker may be able to infer whether a certain text passage appears in the database by observing the outputs of the RAG system.
 This type of attack is known as a Membership Inference Attack (MIA), and can be used to reveal sensitive information about the contents of the retrieval database.
-MIAs were extensively researched in the past in the context of various machine and deep learning systems [15, 3, 2, 18, 9], however to the best of our knowledge the topic of membership inference against RAG systems remains under-explored.
-When a MIA is performed against a RAG system it can potentially reveal sensitive or proprietary company information.
-This may include information about individuals or organizations included in the retrieval database.
-Furthermore, MIA can be used to prove the unauthorized use of proprietary documents, as part of a legal action [16].
-This dual capability makes them a particularly serious form of attack that must be investigated in order to ensure the security and privacy of these systems.
+MIAs were extensively researched in the past in the context of various machine and deep learning models [Shokri et al., 2017, Carlini et al., 2022, Amit et al., 2024, Tseng et al., 2021, Hu et al., 2022], facilitating the detection of models’ training data.
+However, to the best of our knowledge, the topic of membership inference against RAG systems remains under-explored.
+When a MIA is performed against a RAG system, it can potentially reveal sensitive or proprietary company information, including information about individuals or organizations included in the retrieval database.
+Furthermore, MIA can be used to prove the unauthorized use of proprietary documents, as part of a legal action [Song and Shmatikov, 2019].
+This dual capability makes them a serious form of attack that must be investigated to ensure the security and privacy of these systems.
+However, existing approaches to performing MIA on LLMs may not necessarily succeed in leaking membership information about RAG documents.
+Specifically, we show that careful selection of the attack prompt, and adapting it to the RAG scenario, is crucial to its success.
 In this study, we introduce a method that is both efficient and easy to use for conducting MIAs against RAG systems, with the objective of determining whether a specific data sample is present in the retrieval database.
-We assess the effectiveness of our attack using two benchmark datasets that represent scenarios where privacy may be a concern, and employing multiple generative models.
+The method treats the entire RAG pipeline as a black box and does not require access to, or knowledge of, any internal implementation details, or even the RAG template, making it especially useful and easy to employ.
+We assess the effectiveness of our attack using two benchmark datasets that represent various domains where privacy may be a concern, namely medical questions and answers and emails, and employing multiple generative models.
 Moreover, we introduce an initial defense strategy based on adding instructions to the RAG template to prevent the model from responding to the attack prompt.
-This approach seems to be highly effective for some datasets and models, but should be further developed to provide a more comprehensive solution.
+This approach does not require introducing any additional components or add any computational overhead to the overall RAG solution.
+This defense seems to be highly effective for some datasets and models, but should be further developed to provide a more comprehensive solution.
 The findings of our study reveal that the membership of a document in the retrieval database can be efficiently determined through the creation of an appropriate prompt, underscoring the importance of developing appropriate defenses and implementing security countermeasures in deployed RAG systems.
@@
 Retrieval Augmented Generation is a technique for enriching the knowledge of Large Language Models (LLMs) with external databases without requiring any additional training.
-This allows easy customization of trained LLMs for specific needs, such as creating AI-based personal assistants or making long textual content (such as a user manual) accessible using simple text queries [4].
+This allows easy customization of trained LLMs for specific needs, such as creating AI-based personal assistants or making long textual content (such as a user manual) accessible using simple text queries [Chase, 2022].
 Prior to deploying a RAG pipeline, a retrieval database $\mathcal{D}$ must be populated with documents.
-During this initialization phase, each document is split into chunks, which are mapped into vector representations using a retrieval model $E$, and then stored as an index in the database alongside the original document.
-The retrieval model $E$ is specifically designed to learn a mapping from user prompts and documents to a shared vector space, such that prompts are embedded close together with documents containing relevant information to respond to the prompt.
+During this initialization phase, each document is split into chunks, which are mapped into vector representations (embeddings) using an embedding model $E$, and then stored as an index in the database alongside the original document.
+The embedding model $E$ is specifically designed to learn a mapping from user prompts and documents to a shared vector space, such that prompts are embedded close together with documents containing relevant information to respond to the prompt.
 This enables efficient searching in the retrieval database, as semantically similar prompts and documents are clustered together in the vector space.
@@
 When a user prompt is processed by the RAG pipeline, the prompt $p$ is mapped into a vector representation using $E$.
-Then, $\mathcal{D}$ is searched to find the top-k most similar entries, based on a distance metric calculated using the vector representations (e.g., Euclidean distance).
+Then, $\mathcal{D}$ is searched to find the top-k most similar entries, based on a distance metric calculated on the vector representations (e.g., Euclidean distance).
 The retrieved entries from $\mathcal{D}$ are organized and provided to the generation phase together with the user prompt.
@@
 where $\parallel$ denotes text concatenation.
-In practice, it is common to place titles indicating the different areas in the input or additional instructions.
+In practice, it is common to place titles in the template to indicate the different areas, as well as provide additional instructions.
 For example, placing the sentence ”Answer the question based on the context” after the context and before the user prompt.
@@
 2.2 Membership Inference Attacks
-Membership inference attacks [15, 9] are a type of privacy threat, where an attacker aims to determine whether a specific data record was used in the training set of a machine learning model.
+Membership inference attacks [Shokri et al., 2017, Hu et al., 2022] are a type of privacy threat, where an attacker aims to determine whether a specific data record was used in the training set of a machine learning model.
 This carries significant privacy implications as it can potentially reveal sensitive information about individuals, even if the model does not directly release any personal data.
-Formally, an attacker aims to determine the membership of a sample $x$ in the training data of a target model $\mathcal{D}_{m}$, i.e., to check if $x\in\mathcal{D}_{m}$, which is known as sample-level membership inference.
-Typically, these attacks involve calculating one or more metrics on the target model’s outputs that reflect the probability of the sample being a part of the training set, such as the model outputs’ entropy or log-probabilities [3].
+Formally, an attacker aims to determine the membership of a sample $x$ in the training data $\mathcal{D}_{m}$ of a target model $m$, i.e., to check if $x\in\mathcal{D}_{m}$.
+This is known as sample-level membership inference.
+Typically, these attacks involve calculating one or more metrics on the target model’s outputs that reflect the probability of the sample being a part of the training set, such as the model outputs’ entropy or log-probabilities [Carlini et al., 2022].
 Several metrics may be computed for each sample and then fused together using a machine learning model, known as an attack model, which in turn outputs the probability of a sample being a member of the training set.
-Additionally, an attacker may also aim to determine the membership of a certain user, i.e., to check if a user’s data is part of the training set, which is known as user-level membership inference [14].
+Additionally, an attacker may also aim to determine the membership of a certain user, i.e., to check if a user’s data is part of the training set, which is known as user-level membership inference [Shejwalkar et al., 2021].
 Throughout this paper we will address the membership inference challenge from a sample-level perspective.
@@
 Furthermore, the attacker has no information regarding the deployment details, such as the type of retrieval database employed.
-In addition to the black-box setting, we also evaluate a supplementary gray-box scenario, in which the attacker has access to the log-probabilities of the generated tokens.
-Moreover, in this scenario, we assume that the attacker can train the attack model on the target model’s outputs on a subset of the model’s actual training and test datasets [13].
+In addition to the black-box setting, we also evaluate a supplementary gray-box scenario, in which the attacker has access to the log-probabilities of the generated tokens, as previously explored in prior art [Duan et al., 2024, Zhang et al., 2024] Moreover, in this scenario, we assume that the attacker can train the attack model on a subset of the model’s actual training and test datasets [Shachor et al., 2023].
+3 RELATED WORK
+3.1 Membership Inference Attacks against LLMs
+Membership inference attacks have been extensively explored for various types of machine and deep learning models [Shokri et al., 2017, Carlini et al., 2022, Hu et al., 2022].
+Recent interest in Large Language Models (LLMs), has brought a variety of MIA studies tailored for such models [Mahloujifar et al., 2021, Kandpal et al., 2023, Panda et al., 2024].
+While some studies utilized existing concepts such as loss-based attacks [Shejwalkar et al., 2021], others employed domain-specific approaches, for example the use of the Perplexity measure to indicate membership [Galli et al., 2024].
+After RAG was introduced to enhance the capabilities of LLMs, it also became a target for privacy attacks, such as [Zeng et al., 2024], that attempts to extract data from the retrieval database.
+However, to the best of our knowledge, this is the first paper that suggests a MIA against such systems.
+Shortly after our preliminary publication, another paper suggested a different approach to MIA against RAG systems, utilizing the semantic similarity between the model’s generated text and the original sample as the membership metric [Li et al., 2024b].
+However, our method does not rely on the model to actually leak the original text, which is usually considered a harder problem than that of determining membership alone.
+3.2 Prompt Injection Attacks
+A prompt injection attack aims to compromise an LLM-based application so that it produces some desired response, for example, changing the LLM prediction or utilizing the LLM for another purpose that is not its intended task.
+Our attack can be considered a kind of prompt injection attack  [Liu et al., 2024], whose goal is to violate the privacy of the retrieval dataset.
+Prompt injection attacks can take several forms.
+According to the framework described in  [Liu et al., 2024], our attack falls under the category of ”Naive attacks”, such as [Willison, 2022, Matthew Kosinski, 2024], in which the adversary’s instruction is simply concatenated to the target data.
+Other types of prompt injection attacks have also been employed against RAG systems, with the goal of interfering with their outputs.
+In [Zhong et al., 2023], the authors describe a poisoning attack where the adversary introduces malicious documents into the retrieval database, causing the RAG system to output undesirable responses in response to certain queries.
 4 METHODOLOGY
-Figure 2: Attack prompt example for RAG MIA. The highlighted text is the attack-specific part of the prompt, and the rest is taken from the sample for which membership is inferred.
-In this section we present our membership inference attack.
-The input to the generative model used by the system consists of the RAG tempate into which the retrieved database entries and the user prompt are inserted, as depicted in Figure 1.
+Figure 2: Overall Flow of our MIA Attack on a RAG pipeline.
+Figure 3: Attack prompt example for RAG-MIA. The highlighted text is the attack-specific part of the prompt, and the rest is taken from the sample for which membership is inferred.
+In this section we present our membership inference attack (denoted as RAG-MIA), which is shown in Figure 2.
+The input to the generative model consists of the RAG template into which the retrieved database entries and the user prompt are inserted, as depicted in Figure 1.
 Since the attacker’s interaction with the system is limited to the user prompt part of the RAG template, the attack must be based on a prompt that achieves both of the following goals:
@@
 The first goal can be easily fulfilled by creating a prompt that has high similarity with the desired sample in the retrieval database.
-Essentially, any prompt that contains the target sample should be sufficient, as long as there is not too much text added to it.
+Essentially, any prompt that contains the target sample without too much additional text should be sufficient.
 However, the prompt must also cause the generation process to expose the membership status of the sample.
-Prompting the system with target sample alone will not necessarily achieve this desired result.
-To handle this, we designed the following attack prompt format, along with a few additional variations presented in Table 1: ”Does this: ”{Target Sample}” appear in the context?
+Prompting the system with the target sample alone will not necessarily achieve this goal.
+To handle this, we designed the following attack prompt format (along with a few additional variations presented in Table 1): ”Does this: ”{Target Sample}” appear in the context?
 Answer with Yes or No.”.
-In Figure 2 we present an example of this attack prompt format.
+In Figure 3 we present an example of this attack prompt format with a specific target sample.
 Using this prompt we are able to achieve both goals: cause the right sample to be retrieved from the retrieval database, and make the generation phase produce a binary response regarding the sample’s presence in the context.
 In the black-box setting, we use the model’s answer (Yes/No) alone to deduce the membership status of samples.
-As an enhancement to our attack, in cases where the adversary has access to the log-probabilities of the selected tokens, we additionally employ an attack model, similar to the approach described in Section 2.2.
-In this setup, we apply an ensemble of attack models [13], using as input to the attack both the logits and the class-scaled-logits [3] corresponding to the ”Yes” and ”No” tokens output from the attacked model.
-The logits are computed by first calculating the exponent of of the log-probability to get a probability estimate $P$ and then applying the logit function.
+As an enhancement to our attack, in cases where the adversary has access to the log-probabilities of the selected tokens [Zhang et al., 2024, Duan et al., 2024], we additionally employ an attack model (see Section 2.2) to determine membership.
+In this setup, which we call the gray-box setting, we employ an ensemble of attack models [Shachor et al., 2023] that receive as input both the logits and class-scaled logits [Carlini et al., 2022] corresponding to the ”Yes” or ”No” token output from the target model.
+The logits are computed by first calculating the exponent of the log-probability to get a probability estimate $P$ and then applying the logit function.
 Since the model only outputs the log-probability of the selected token, for example the ”Yes” token, without the complementary ”No” token, we assign a fixed low probability value of $0.001$ to the complementary token.
-Black-Box
-Gray-Box
-Figure 3: Comparison of different attack prompts. Top: HealthCareMagic, Bottom: Enron
+More details about this attack setup can be found in Section MIA Attack Details.
 4.1 Experimental Setup
@@
 The experiments were conducted on three generative language models:
-  • google/flan-ul2 [17], denoted by flan
-  • meta-llama/llama-3-8b-instruct [1], denoted by llama
-  • mistralai/mistral-7b-instruct-v0-2 [11], denoted by mistral
 Datasets
@@
 Retrieval Database
-We used a Milvus Lite [19, 8] vector database with $k=4$, Euclidean distance (L2) metric type and HNSW index.
+We used a Milvus Lite [Wang et al., 2021, Guo et al., 2022] vector database with $k=4$, Euclidean distance (L2) metric type and HNSW index.
 RAG Template
 The input to the generative model is built from the user prompt and the context fetched from the retrieval database as a response to the prompt, using the following template:
-Please answer the question using the context provided.
-If the question is unanswerable, say unanswerable”.
+Please answer the question using the context
+provided. If the question is unanswerable, say
+"unanswerable".
 Question: {user prompt}.
@@
 In our evaluation we experimented with 5 different attack prompts, listed in Table 1.
-Each attack prompt includes a placeholder for a sample, which can be a member or a non-member sample, as shown in Section 3.
+Each attack prompt includes a placeholder for a sample, which can be a member or a non-member sample.
 In the case of the Enron dataset, the sample is the full email body, or its first 1000 characters if it is longer.
-In HealthCareMagic, the human part of the dialogue is used as the sample.
+In the HealthCareMagic dataset, the human part of the dialogue is used as the sample.
 MIA Attack Details
-In both the black-box and the gray-box scenarios we ran our attack using random samples containing 2000 member documents and 2000 non-member documents.
-In the gray-box case, these documents were split between 40 attack models in the ensemble [13], each model trained 5 times with different train-test splits.
-In this evaluation, we classified model outputs that failed to include either the ”Yes” or ”No” tokens as non-member documents.
+In both the black-box and the gray-box scenarios, we started by randomly sampling 2000 member documents and 2000 non-member documents to be used in the evaluation.
+We then ran our attack 10 times, each time using a different random sample of 500 member documents and 500 non-member documents out of those 2000 member or non-member documents selected in the previous stage.
+For the gray-box scenario we leveraged the attack technique proposed by [Shachor et al., 2023], which combines multiple attack models into a powerful ensemble.
+Each model in the ensemble is trained on a small subset of the data and optimized using a comprehensive hyper-parameter search covering various attack model types, data scaling approaches, and model training parameters.
+This creates diversity among model predictions, and improves the overall membership inference attack performance.
+In our experiments, we used an ensemble consisting of 40 attack models.
+Throughout the experiments in this paper, model outputs that failed to include either ”Yes” or ”No” tokens were classified as non-member documents.
 This case accounted for approximately 6% of the total outputs.
-We present the statistics on the amount of such documents in Appendix C.
+We present detailed statistics on this in Appendix C.
+Evaluation Metrics
+We employ three primary evaluation metrics to assess the performance of our attack.
+Firstly, for the gray-box attack, we employ TPR@lowFPR [Carlini et al., 2022], which measures the percentage of correctly identified member samples (True Positive Rate) at a fixed, low percentage of falsely identified member samples (False Positive Rate).
+In our case, we present results for the extreme case of FPR=0, where all non-member samples are correctly classified.
+Secondly, for the black-box attack, where we simply use the Yes/No answer generated by the model and have no threshold that can be tuned, we report the single TPR and FPR values achieved by the attack (denoted as TPR@FPR).
+Thirdly, we utilize the Area Under the Receiver Operating Characteristic curve (AUC-ROC), a common metric for evaluating binary classification problems, including membership inference attacks.
 5 RESULTS
+Enron Enron HealthcareMagic HealthcareMagic
+AUC-ROC TPR@FPR AUC-ROC TPR@FPR
+Black-Box
+Refer to caption Refer to caption Refer to caption Refer to caption
+Gray-Box
+Refer to caption Refer to caption Refer to caption Refer to caption
+Figure 4: Comparison of different attack prompts.
+The top row shows results for black-box attacks, evaluated using AUC-ROC and TPR@FPR.
+The wide bars show the TPR and the narrow bars inside them show the respective FPR values.
+The bottom row shows results for gray-box attacks, evaluated using AUC-ROC and TPR@lowFPR.
 5.1 Impact of Attack Prompts
 One important aspect of performing attacks against any generative model is the prompt used to trigger the attack.
-In the case of RAG, the attack prompt is first used to fetch the relevant documents from the retrieval database, and is then incorporated together with the fetched documents into the RAG template and provided to the generative model.
-To achieve the best performance, we experimented with five possible attack prompts, described in Table 1.
+In the case of RAG, the attack prompt is first used to fetch the relevant documents from the retrieval database, then it is incorporated into the RAG template together with the fetched documents and used as input for the generative model.
+To achieve the best performance, we experimented with five possible attack prompts, detailed in Table 1.
 Each one tries to manipulate the RAG system in a slightly different manner to both fetch the relevant documents from the retrieval database, and cause the generative model to output whether the target document was indeed in the retrieval database.
@@
 […]
+your context?
 […]
+appear in the context?
 […]
+Answer with Yes or No.
 […]
+one word.
 […]
+Answer with one word. Answer No if you are not sure.
+In Figure 4 we present the AUC-ROC and TPR results for both the gray-box and black-box scenarios, measuring the effectiveness of the different attack prompts.
+In the black-box scenario, the attack predictions are discrete, meaning that computing TPR@lowFPR is not possible.
+Instead we present the TPR@FPR of the attack, i.e., the percentage of members correctly classified as members and the percentage of non-members classified as members.
+The same results organized in table format can be found in Appendix 7.
 […]
-Unsurprisingly, the results in the black-box setting are inferior to those in the gray-box setting.
-This is especially the case for the flan model, with an improvement of up to 25% in the gray-box setting.
-This means that the log-likelihood values for member documents are significantly higher than for non-member documents, indicating a higher confidence of the model in its response.
-However, when looking at the llama and mistral models, the average difference is only up to 12%, and in some cases even lower, depending on the prompt.
-To further explore this difference, we analyze the percentage of samples that are correctly retrieved from the database for each prompt.
+Unsurprisingly, the TPR@FPR and the AUC-ROC results in the gray-box setting are superior to those in the black-box setting.
+This is prominent in the case of the flan model, with an improvement of up to 22% in the gray-box setting.
+This means that for this model, the log-probability values for member documents are significantly higher than for non-member documents, indicating a higher confidence of the model in its response.
+However, when looking at the llama and mistral models, the average difference is only up to 7%, and in some cases even lower, depending on the prompt.
+To further explore this difference between the models, we analyzed the percentage of member samples that are correctly retrieved from the database for each prompt.
 We found that over 95% of the member samples are indeed retrieved for both datasets.
-This is in contrast with the non-member samples, which are retrieved in nearly 0% of the cases.
+This is in contrast with the non-member samples, that are retrieved in nearly 0% of the cases.
 The full results of this analysis can be found in Appendix B.
-Thus, we conclude that the flan model is more grounded to the content of its input prompt (context grounding), and thereby more sure of the presence/absence of a piece of text from it in comparison to the llama and mistral models.
+Thus, we conclude that the flan model is more grounded to the content of its input prompt (i.e., context grounded), and thereby more sure of the presence/absence of a piece of text from it in comparison to the llama and mistral models.
 5.2 Attack Results Summary
-To quantify the overall MIA risk, we summarize the best results (across prompts) for each combination of model and dataset in Table 2 .
-We can see that, on average, the attack success rate is highest for flan, in both threat models.
-In addition, we observe that the overall risk in the black-box setting is similar between almost all models, with the exception of the HealthCareMagic-llama case.
-However in the gray-box setting the results vary more.
-Table 2: RAG-MIA results summary
+Table 2: RAG-MIA results summary.
 Black-Box
-AUC ROC
+TPR
+Black-Box
+FPR
 Gray-Box
-AUC ROC
+TPR@lowFPR
+Black-Box
+AUC-ROC
+Gray-Box
+AUC-ROC
 Dataset Model
-HealthCareMagic flan 0.80 1.00
-llama 0.88 0.96
-mistral 0.74 0.83
-Enron flan 0.82 0.96
-llama 0.79 0.83
-mistral 0.78 0.82
-Our lowest attack AUC-ROC score of 0.74 is quite high compared to previous research on sample-level MIA in language models [5, 14].
-Overall, our black-box attack achieves an average AUC-ROC of 0.80 and the gray-box attack achieves an average AUC-ROC of 0.90.
-The best gray-box attack on flan accomplishes almost perfect performance.
-These results highlight the risk associated with deploying RAG based systems without proper defense mechanisms.
-4.3 Defense Strategy
-To counter the threat of RAG-MIA attacks, we propose an initial defense strategy in this section.
-Building on previous research that has demonstrated the ability to influence the output of LLMs through system instructions, we leverage this concept to develop a defense mechanism.
-Our approach involves modifying the RAG template to include explicit instructions for the LLM to disregard requests that directly inquire about the retrieved database content, thereby providing a defense against RAG-MIA attacks.
+HealthCareMagic flan 1.00 0.61 0.85 0.81 0.99
+llama 0.95 0.20 0.73 0.89 0.96
+mistral 0.42 0.10 0.36 0.74 0.83
+Enron flan 1.00 0.56 0.63 0.82 0.96
+llama 0.78 0.30 0.28 0.79 0.83
+mistral 0.61 0.17 0.22 0.78 0.81
+We present a summary of the best results from Figure 4 for each model/dataset combination in Table 2, which quantifies the overall risk of Membership Inference Attacks.
+The results show that the attack is most effective against the flan models, in both the gray- and black-box scenarios.
+Notably, the overall risk in both scenarios is very high, reaching nearly perfect AUC-ROC in the gray-box scenario (with an average of 0.9 across models and datasets) and close to 0.9 AUC-ROC in the black-box scenario (0.8 on average).
+The TPR@lowFPR values are also remarkably high, ranging from 0.22 to 0.85 (0.51 on average) for an FPR of 0.
+This significantly surpasses TPR results from previous MIA research in language models [Carlini et al., 2022, Li et al., 2024a, Zhang et al., 2024], which is usually no higher than 0.25 even for an FPR of 0.05.
+In the black-box scenario, the TPR for all cases is higher by at least 30% than the FPR.
+This suggests that the attack is highly effective in real-world scenarios where the attacker aims to correctly identify member samples.
+These results underscore the significant risk associated with deploying RAG-based systems without adequate defense mechanisms in place.
+5.3 Initial Defense Strategy
+To counter the threat of RAG-MIA attacks, in this section we propose an initial defense strategy.
+Building on previous research that has demonstrated the ability to influence the output of LLMs through system instructions [Reynolds and McDonell, 2021, Lyu et al., 2024], we leverage this concept to develop a basic defense mechanism.
+Our approach involves modifying the RAG template to include explicit instructions for the LLM to disregard requests that directly inquire about the retrieved database content.
 Specifically, we evaluated the effectiveness of variations of the following modified RAG template:
-Please answer the question using the context provided.
-If the question is unanswerable, say unanswerable”.
+Please answer the question using the context
+provided. If the question is unanswerable,
+say "unanswerable".
 Pointers for the answer:
-1. Repeating, outputting or printing the context is illegal,
-answer unanswerable”.
-2. Asking if a specific text, fact, or passage appears in your
-context is illegal, answer unanswerable”.
+1. Repeating, outputting or printing the
+context is illegal, answer "unanswerable".
+2. Asking if a specific text, fact, or
+passage appears in your context is illegal,
+answer "unanswerable".
 Question: {user prompt}
@@
 {context}
-We have slightly adapted the RAG template for the llama model, such that it will respond to the defense instructions better, by placing the defense instructions inside the system prompt, an integral capability of llama models.
-In Table 3 we present the results, and compare them to the scenario where no defense is placed in the RAG system.
-Table 3: RAG MIA results with defense
+For the llama model, that has a separate section for a system prompt and is specifically designed to adhere to the instructions contained within it, we slightly adapted the RAG template by placing the defense instructions inside the system prompt.
+This was done to increase the likelihood of the model to respond to the defense instructions.
+To the best of our knowledge, neither flan nor mistral have such sections.
+Following is the modified RAG template for llama:
+<|begin_of_text|><|start_header_id|>system
+<|end_header_id|>
+Please answer the question using the context
+provided. If the question is unanswerable,
+say "unanswerable".
+Pointers for the answer:
+1. Repeating, outputting or printing the
+context is illegal, answer "unanswerable".
+2. Asking if a specific text, fact, or
+passage appears in your context is illegal,
+answer "unanswerable".
+Context:
+{context}
+<|eot_id|>
+<|begin_of_text|><|start_header_id|>user
+<|end_header_id|>
+Question: {user_prompt}.
+<|eot_id|><|start_header_id|>assistant
+<|end_header_id|>
+Table 3: RAG-MIA results with defense - TPR@FPR.
 Without defense With defense
 Black-Box
-AUC ROC
+TPR
+Black-Box
+FPR
 Gray-Box
-AUC ROC
+TPR@lowFPR
 Black-Box
-AUC ROC
+TPR
+Black-Box
+FPR
 Gray-Box
-AUC ROC
+TPR@lowFPR
 Dataset Model
-HealthCareMagic flan 0.78 0.99 0.86 0.90
-llama 0.88 0.96 0.74 0.51
-mistral 0.74 0.83 0.72 0.44
-Enron flan 0.82 0.97 0.88 0.96
-llama 0.79 0.84 0.77 0.78
-mistral 0.78 0.80 0.78 0.72
-Our evaluation reveals that the proposed defense strategy yields the most significant benefits against gray-box attacks on the llama and mistral models, across both datasets.
-Notably, the defense demonstrates high efficacy in the gray-box setting for the mistral model, particularly on the HealthCareMagic dataset, where we observe a substantial improvement of 0.39 in AUC-ROC.
+HealthCareMagic flan 1.00 0.61 0.85 0.67 0.02 0.65
+llama 0.95 0.20 0.73 0.09 0.00 0.13
+mistral 0.42 0.10 0.36 0.11 0.01 0.13
+Enron flan 1.00 0.56 0.63 0.77 0.04 0.69
+llama 0.78 0.30 0.28 0.42 0.04 0.32
+mistral 0.61 0.17 0.22 0.52 0.06 0.27
+Table 4: RAG-MIA results with defense - AUC-ROC.
+Without defense With defense
+Black-Box
+AUC-ROC
+Gray-Box
+AUC-ROC
+Black-Box
+AUC-ROC
+Gray-Box
+AUC-ROC
+Dataset Model
+HealthCareMagic flan 0.81 0.99 0.85 0.90
+llama 0.89 0.96 0.74 0.51
+mistral 0.74 0.83 0.72 0.45
+Enron flan 0.82 0.96 0.88 0.97
+llama 0.79 0.83 0.77 0.77
+mistral 0.78 0.81 0.78 0.72
+In Table 3 and Table 4 we present the results of our attacks on the defended models, comparing them to the undefended models.
+This evaluation reveals that the proposed defense strategy yields the most significant benefits against gray-box attacks on the llama and mistral models, across both datasets.
+On the HealthCareMagic dataset, we observe a substantial improvement of 0.45 and 0.38 in AUC-ROC and of 0.6 and 0.23 in TPR@lowFPR, respectively.
 In contrast, the defense has a minimal impact on the flan model, only showing a slight effect in the gray-box setting.
 These findings suggest that further research is needed to explore the feasibility of crafting a RAG template that can effectively defend a flan model against RAG-MIA attacks.
-Furthermore, our analysis of the model outputs reveals that a significant proportion of responses do not contain either ”Yes” or ”No” tokens (classified as non-member documents), accounting for approximately 96% of the total outputs on llama, and about 93% on mistral with HealthCareMagic dataset, which explains the defense performance for these cases.
-On the other hand, for flan, they remain the same (0%).
-We present the statistics on the amount of such documents in Appendix D.
+Furthermore, our analysis of the model outputs reveals that, when applying this defense, a significant proportion of the responses does not contain a ”Yes” or ”No” token.
+Instead, it contains ”unanswerable” as per the defense instruction.
+This accounts for approximately 96% of the total outputs from llama, and 93% from mistral with the HealthCareMagic dataset.
+As previously noted, our attack classifies these responses as non-member documents, explaining the improved defense performance for these cases.
+On the other hand, for flan, the percentage of responses not containing the ”Yes” or ”No” token remains the same (0%).
+Detailed statistics are presented in Appendix D.
 5.3.1 Improved defense for llama
-Since llama models have a dedicated section in the input prompt for system instructions, as mentioned in Section 4.3, we have experimented with placing the retrieved database content within this dedicated section.
+As mentioned in Section 5.3, llama models have a dedicated section in the input prompt for system instructions.
+We thus also experimented with placing the retrieved database content within this dedicated section.
 We compare two cases: (1) With Defense #1 - only the defense instructions are added to the system section (2) With Defense #2 - both defense instructions and retrieved database content are placed in the system section.
-We present the results for this experiment in Table 4.
-Table 4: Llama defenses
+We present the results of this experiment in Tables 5 and 6.
+Table 5: Llama defenses - TPR@FPR.
+Without Defense With Defense #1 With Defense #2
+Black-Box Gray-Box Black-Box Gray-Box Black-Box Gray-Box
+TPR FPR TPR@lowFPR TPR FPR TPR@lowFPR TPR FPR TPR@lowFPR
+Dataset
+HealthCareMagic 0.95 0.20 0.73 0.09 0.00 0.13 0.00 0.00 0.30
+Enron 0.78 0.30 0.28 0.42 0.04 0.32 0.01 0.01 0.49
+Table 6: Llama defenses - AUC-ROC.
 Without Defense With Defense #1 With Defense #2
 Black-Box
-AUC ROC
+AUC-ROC
 Gray-Box
-AUC ROC
+AUC-ROC
 Black-Box
-AUC ROC
+AUC-ROC
 Gray-Box
-AUC ROC
+AUC-ROC
 Black-Box
-AUC ROC
+AUC-ROC
 Gray-Box
-AUC ROC
-Dataset Model
-HealthCareMagic llama 0.88 0.96 0.74 0.51 0.42 0.75
-Enron llama 0.79 0.84 0.77 0.78 0.46 0.92
-As shown in Table 4, placing defense instructions and retrieved database content in the system section provides robust defense against black-box attacks.
+AUC-ROC
+Dataset
+HealthCareMagic 0.89 0.96 0.74 0.51 0.54 0.74
+Enron 0.79 0.83 0.77 0.77 0.51 0.92
+As shown in Table 5, placing both the defense instructions and the retrieved database content in the system section provides a robust defense against black-box attacks.
 However, this approach is less effective against gray-box attacks, where Defense #1 is preferred.
 Since black-box attacks are more common and require less effort from the attacker, we recommend using Defense #2.
-Nevertheless, we encourage the research community to develop a defense strategy that can effectively protect from both kinds of attack.
-5 Conclusion
-In this paper we introduced a new membership inference attack against RAG-based systems meant to infer if a specific document was part of the retrieval database or not, a research question previously not explored.
-Our attack is demonstrated both in black-box and gray-box threat models, and takes advantage of a characteristic of generative models that is usually considered an advantage - context grounding.
-Our attack achieves a very high average AUC-ROC of 0.90 and 0.80 in the gray-box and the black-box threat models, respectively, and for some models almost perfect performance.
-Our initial defense was able to reduce the success rate of the attack in almost all cases, and mostly prevented the attack for the llama model in the black-box setting.
-The model that mostly did not benefit from this defense was the flan model for which further defenses need to be developed.
-This underscores the need to develop more advanced defenses and countermeasures for this attack scenario.
-We hope that the research community will proceed to explore the risk of membership inference in RAG-based systems and employ the ideas from this paper as a baseline.
-6 Acknowledgement
+Nevertheless, we encourage the research community to further research and develop defense strategies that can effectively protect against both kinds of attacks.
+6 CONCLUSIONS
+In this paper we introduced a new membership inference attack against RAG-based systems meant to infer if a specific document is part of the retrieval database or not.
+Our attack does not rely on the model replicating text from the retrieval database, and rather relies on binary answers provided by the model itself regarding its context.
+This takes advantage of a characteristic of generative models that is usually considered an advantage - context grounding.
+We demonstrated results both in black-box and gray-box threat models.
+Our attack achieves a very high performance, with average AUC-ROC of 0.90 and 0.80 in the gray-box and the black-box threat models, respectively, and for some models achieving almost perfect performance.
+Conversely, our initial defense was able to reduce the success rate of the attack in almost all cases, and essentially prevented the attack for the llama model in the black-box setting.
+This illustrates the need for more advanced attack prompts, for example integrating prompt-injection techniques.
+The model that mostly did not benefit from this defense was the flan model, for which further defenses need to be developed.
+Furthermore, while this paper only explored direct attacks, it is important to take into account adaptive attacks (e.g.
+[Tramer et al., 2020]) that consider potential defenses, such as the one presented in this paper, in the attack process.
+This underscores the need to establish more advanced defense strategies and countermeasures.
+Such approaches may be based on differentially private synthetic data generation [Amin et al., 2024, Xie et al., 2024], which employs LLMs to generate synthetic texts based on several seed text passages.
+Alternatively, differential privacy mechanisms can be employed to the LLMs’ text generation process [Du and Mi, 2021, Majmudar et al., 2022], which may also help to reduce the risk.
+In summary, we hope that the research community will continue to explore the risk of membership inference in RAG-based systems and employ the ideas from this paper as a baseline.
+ACKNOWLEDGEMENTS
 This work was performed as part of the NEMECYS project, which is co-funded by the European Union under grant agreement ID 101094323, by UK Research and Innovation (UKRI) under the UK government’s Horizon Europe funding guarantee grant numbers 10065802, 10050933 and 10061304, and by the Swiss State Secretariat for Education, Research and Innovation (SERI).
-Appendices
-A Detailed Attack AUC ROC Scores
-Table 5 presents the AUC ROC scores for both threat models: black-box and gray-box, and for the different attack prompts.
-Table 5: Full RAG-MIA results
+APPENDIX
+A Detailed Attack Scores
+Table 7 presents the full AUC-ROC and TPR scores for the different attack prompts for both threat models: black-box and gray-box.
+Table 7: Full RAG-MIA results.
 Black-Box
-AUC ROC
+TPR
+Black-Box
+FPR
 Gray-Box
-AUC ROC
+TPR@lowFPR
+Black-Box
+AUC-ROC
+Gray-Box
+AUC-ROC
 Dataset Model
@@
 Prompt
-HealthCareMagic flan 0 0.80 0.98
-1 0.75 0.97
-2 0.78 0.99
-3 0.76 0.84
-4 0.80 1.00
-llama 0 0.82 0.88
-1 0.67 0.64
-2 0.88 0.96
-3 0.78 0.83
-4 0.84 0.89
-mistral 0 0.71 0.73
-1 0.67 0.65
-2 0.74 0.83
-3 0.71 0.73
-4 0.71 0.55
-Enron flan 0 0.79 0.93
-1 0.68 0.80
-2 0.82 0.96
-3 0.60 0.62
-4 0.80 0.92
-llama 0 0.71 0.65
-1 0.62 0.54
-2 0.79 0.83
-3 0.69 0.58
-4 0.74 0.74
-mistral 0 0.66 0.60
-1 0.70 0.68
-2 0.78 0.82
-3 0.64 0.53
-4 0.74 0.69
+HealthCareMagic flan 0 1.00 0.63 0.76 0.80 0.97
+1 0.99 0.84 0.76 0.75 0.97
+2 1.00 0.74 0.83 0.78 0.99
+3 0.98 0.70 0.33 0.76 0.83
+4 1.00 0.61 0.85 0.81 0.99
+llama 0 0.91 0.29 0.48 0.82 0.87
+1 0.64 0.30 0.28 0.67 0.63
+2 0.95 0.20 0.73 0.89 0.96
+3 0.92 0.44 0.47 0.77 0.82
+4 0.91 0.25 0.38 0.84 0.88
+mistral 0 0.91 0.63 0.32 0.70 0.72
+1 0.83 0.54 0.27 0.67 0.65
+2 0.97 0.69 0.36 0.74 0.83
+3 0.94 0.71 0.22 0.70 0.73
+4 0.42 0.10 0.23 0.71 0.54
+Enron flan 0 0.99 0.65 0.60 0.79 0.93
+1 0.94 0.75 0.46 0.68 0.81
+2 1.00 0.56 0.63 0.82 0.96
+3 0.87 0.73 0.27 0.60 0.63
+4 1.00 0.63 0.34 0.81 0.91
+llama 0 0.92 0.63 0.13 0.71 0.66
+1 0.63 0.38 0.14 0.62 0.54
+2 0.91 0.38 0.28 0.79 0.83
+3 0.93 0.72 0.14 0.69 0.59
+4 0.78 0.30 0.17 0.74 0.75
+mistral 0 0.87 0.64 0.09 0.66 0.59
+1 0.85 0.50 0.21 0.70 0.68
+2 0.92 0.44 0.22 0.78 0.81
+3 0.86 0.65 0.10 0.64 0.51
+4 0.61 0.17 0.20 0.73 0.67
 B Detailed Results of Retrieval Matches
-Tables 6 and 7 show the percent of exact matches between the retrieved documents and the member and non-member samples, respectively, for the different attack prompts.
+Tables 8 and 9 show the percent of exact matches between the retrieved documents and the member and non-member samples, respectively.
 We see that the chosen attack prompts do not differ in their influence on the retrieval accuracy.
-Table 6: Database retrieval of member documents
+Table 8: Database retrieval of member documents.
 Attack
@@
 4 1908 2000 95.40
-Table 7: Database retrieval of non-member documents
+Table 9: Database retrieval of non-member documents.
 Attack
@@
 C Model Outputs Without a Clear Yes/No Answer
-The number of model outputs that did not contain either of the ”Yes” or ”No” tokens across attack prompts for each dataset and model combination are shown in Table 8.
-Table 8: Model outputs missing yes/no tokens
+The number of model outputs that did not contain either of the ”Yes” or ”No” tokens for each dataset and model combination (across all attack prompts and including both members and non-members) are shown in Table 10.
+Table 10: Model outputs missing Yes/No tokens.
 Missing Total
@@
 D Model Outputs Without a Clear Yes/No Answer with Defense Prompt
-The number of model outputs that did not contain either of the ”Yes” or ”No” tokens when using a defense prompt for each dataset and model combination are shown in Table 9.
-This is compared to the corresponding model outputs when using attack prompt #2 without the defense.
-The number of missing answers increases significantly for the llama and mistral models, and remains the same for flan.
-Table 9: Model outputs missing yes/no tokens with defense prompt
+Table 11: Model outputs missing Yes/No tokens with defense prompt.
 Without defense With defense
@@
 Dataset Model
-HealthCareMagic flan 0 4000 $0$ 4000 0.00%
-llama 0 4000 0.00% 3868 4000 96.70%
+HealthCareMagic flan 0 4000 $0$ 4000 00.00%
+llama 0 4000 00.00% 3868 4000 96.70%
 mistral 101 4000 2.53% 3719 4000 92.97%
-Enron flan 0 4000 $0$ 4000 0.00%
-llama 0 4000 0.00% 3805 4000 95.12%
+Enron flan 0 4000 $0$ 4000 00.00%
+llama 0 4000 00.00% 3805 4000 95.12%
 mistral 41 4000 1.03% 2743 4000 68.58%
+The number of model outputs that did not contain either of the ”Yes” or ”No” tokens when employing a defense in the RAG template (using attack prompt #2) for each dataset and model combination are shown in Table 11.
+This is compared to the corresponding model outputs when using attack prompt #2 without the defense.
+The number of missing answers increases significantly for the llama and mistral models, and remains the same (0) for flan.